A monopolist's demand function is P = 1624 - 4Q, and its total cost function is
(5 marks)
TC = 22,000 + 24Q -4Q2 + 1/3 Q3, where Q is output produced and sold.
i. At what level of output and sales (Q) and price (P) will total profits be
maximized?
ii. At what level of output and sales (Q) and price (P) will total revenue be
maximized?
iii. At what price (P) should the monopolist shut down?
1
Expert's answer
2021-07-25T16:18:03-0400

a . Total Profits are maximized where MR = MC , and MR = d "TR\over dQ", with TR = P ( Q ) , and MC ="dTC\over dQ" . TR = 1624Q - 4Q 2 , so MR = 1624 - 8 Q. MC = 24 - 8Q + Q 2 . MR = MC is 1624 - 8Q = 24 - 8Q + Q 2 , or 1600 = Q 2 , and Q = 40 . With Q = 40 , P = 1464 .

b . Total Revenue is maximized when MR = 0 , or 1624 - 8Q = 0 , or Q = 203 with P = 203 . c . Shut down would occur whenever price ( P ) is less than average variable cost ( AVC ) , or below P = AVC , or 1624 - 4Q = 24 - 4Q + "1\over3" Q 2 , or 1600 ="1\over3" Q 2 , or Q 2 = 4800 , or Q = 69 ( approximately ) . When Q = 69 , P = 1348 , so any price below 1348 would cause the firm to shut down since it is not covering its variable costs .